A right triangle has perimeter 80 inches. The length of the hypotenuse is 34 inches. What is the length of the two shorter sides?
Answer by flame8855(424)

let a be the length of the hypotenuse
let b & c be the lenth of the other two sides.
a=34
c^2 + b^2= a^2
b^2+c^2= 156
a+b+c=80
solve the three equations :
b+c= 80-34=46
c=46-b
b^2 +(46-b)^2=156
b^2-46b+480=0
b=16 or b = 30
for b = 16
c=80-16-34 = 30
for b=30
c=80-30-34=16
so the length of the two other sides are 16 and 30